Supplemental Operations Report – February 3rd

MDRS Supplemental Operations Report 03-Feb-2020

Name of person filing report: Atila Meszaros

Reason for Report: MDRS Hab systems update

Non-nominal systems:
MDRS Generator is non-operational and still on inspection
Loft water tank connection presents a minor leak

Action taken for non-nominal systems:
Since Jan 21st, we have a 15 kW rental generator connected to our power system, meanwhile, we are working to troubleshoot the MDRS generator.
A male to male connection was bought in order to replace it in the line, but the leak it’s not in the connection, it’s in the pipe itself.

Solar— running nominally, solar throughout the day and generator
backup usually on 12 hours a night.

Diesel Reading – 25%, have refill ordered for the end of the week

Propane Reading – 70%

Ethanol Free Gasoline – 2 gallons.

Water (Outpost tank) – 250 gallons

Water (static tank) – 450 gallons

Water (GreenHab) – 130 gallons

Water (loft) – 50 gallons.

Static to Loft Pump used – yes

Water Meter: n/a

Toilet tank emptied: No

Deimos rover used: Off-campus, but on its way back to MDRS this week.

Hours:

Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: Nominal

Hours:

Beginning charge:

Ending charge:

Currently charging: Yes

Spirit rover used: Nominal

Hours:

Beginning charge:

Ending charge:

Currently charging: Yes

Opportunity rover used: Off-campus, but on its way back to MDRS this week.

Hours:

Beginning charge:

Ending charge:

Currently charging:

Curiosity rover used: Nominal.

Hours:

Beginning charge:

Ending charge:

Currently charging: Yes

Notes and action taken on rovers: Nothing to report.

ATV’s used: (Honda, 300, 350.1, 350.2, 350.3):
Honda:
Oil: OK
Battery: non-operational. Management Team tried to jump it, but the ATV won’t keep running.

300:
Currently in town and non-operational.
350-1:
Oil: OK
Fully operational during maintenance running.

350-2:
Oil: OK
Battery: non-operational
Additional safety issue with the damaged handlebars

350-3:
Oil: OK
Fully operational during maintenance running.

Reason for use: Started and checked for operations

Oil Added? No, all OK.

ATV Fuel Used: 0 Gals

# Hours the ATVs were Used: 1

Notes and action taken on ATVs: Nothing more to report

HabCar running? If used, why, where? Started and checked for operations
All systems checked and major inner cleaning by MDRS Management Team.
Transmission fluid: OK
Coolant: 2/3 – OK
Breaker fluid: Full – OK
Oil: Minimum
Power steering fluid: Full – OK
Washer fluid: replacement part to order.

CrewCar running, If used why, where? Started and checked for operations.
All systems checked and major inner cleaning by MDRS Management Team.
Transmission fluid: OK
Coolant: 3/4 – OK
Breaker fluid: Full – OK
Oil: OK
Power steering fluid: Full – OK
Washer fluid: Full – OK (Refilled)

Summary of the internet: Hab and Astronomy accounts nominal. Outpost account nominal, but low in data.

Summary of suits and radios: 5 radios malfunctioning, the main issue seems to be a battery dysfunction, however, new batteries were added and replaced 6 weeks ago for most of the radios. The problem is still on inspection.

Summary of Hab operations:

Camera’s system
All 4 cameras have been installed in the hab:
Inner cameras:
Science Dome
Green Hab
Outside cameras:
From outside the RAM observing the main porch and the station
From outside the Science Dome observing the observatory
All cameras are working perfectly, and the website is still in process. More news about this soon.

Water heater
Up and running. The propane line was secured and the system is already fully operational.

NEW Kitchen Oven
After a long long time, it has arrived. It was picked up yesterday at Salt Lake City and it was finally installed today.

RAM
The breakers and outside plugs have been finally installed, just in time for the soon arrival of the two Rovers.

Lower deck lights
In media res. The full installation will be finished by tomorrow. If the lighting is adequate and enough, the upper deck lights might be changed with this new model too.

Lower deck wall heater
Moved to the right location, connected and fully operational. A replacement propane hose was bought for the water heater system, but it wasn’t finally needed, and it was necessary for the lower deck wall heater, due to the long distance from the propane line in the new correct location

Upper deck loft water tank connection
A male-male connection for the loft tank was bought to be added to the loop, but the leak was caused by the degraded pipe. The leak’s not critical, but a new pipe will be required in the near future.

Supplies added:
2 packs of mop heads
A handle (and a spare) that was missing from a drawer from the IKEA new furniture.

Summary of Outpost operations: Intern Trailer pipelines fully installed and operational.

Summary of GreenHab operations: Crew 220 did a great job on the
GreenHab. Two small disposals containers (organic and non-organic) have been added. Most of the crops look healthy, there has been a good rate of crew’s harvest and it looks like it has been a great season for the Green Hab.

Summary of ScienceDome operations:
Additional wooden pieces for the "island" in the Science Dome have been picked up.

Summary of RAM operations:
Breaker system upgraded and outer plugs for the rovers are fully operational.
The pressurized air pump has a complete hose system already installed and working.
The oil pump has a new hose

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: First Aid Kit for the EVA room seems to be missing.

Questions, concerns, supplies needed and requests: None